OfficialKIzzy:

Given an exponential function for compounding interest, A(x) = P(1.03)x, what is the rate of change? 0.03% 1.03% 3% 103%

563blackghost:

Compound interest function is: \(\bf{A=P(1+\frac{r}{n})^{nt}}\) this means that the decimal of the rate is added with one. \(\bf{1.03 - 1 = .03}\) \(\bf{.03 \rightarrow 3~percent}\)